# Exploring Imbalanced Annotations for Effective In-Context Learning

### v1

@@ A.3 Bayesian Optimizing Conditional Bias
$\mathcal{F}_{\theta}\left(\beta|\mathcal{D}_{r},\mathcal{D}_{b},\bm{w}\right)=\mathcal{N}\left(\mu(\beta),\sigma^{2}(\beta)\right),$
@@
where $\mathcal{N}$ is the Gaussian distribution with a mean function $\mu(\beta)$ and a covariance function $\sigma^{2}(\beta)$. For a data point $\beta_{j}$ sampled from the Gaussian process $\mathcal{N}$, we compute its corresponding function values $\mathcal{F}_{\theta}\left(\beta_{j}|\mathcal{D}_{r},\mathcal{D}_{b},w\right)$.
@@
Acquisition Function. We use the Expected Improvement (EI) criterion to select the point $\beta_{j+1}$ with the maximum expected improvement as the next query point:
@@
$\operatorname{EI}(\beta)$ =$=\left\{\begin{array}[]{ll}\left(\mu_{j}(\beta)-\mathcal{F}_{\theta}\left(\beta^{+}|\mathcal{D}_{r},\mathcal{D}_{b},w\right)-\epsilon\right)\Phi(Z)+\sigma_{j}(\beta)\phi(Z)&\text{ if }\sigma_{j}(\beta)>0,\\ 0&\text{ if }\sigma_{j}(\beta)=0.\end{array}\right.$ (13)
@@
$Z$ $=\frac{\mu_{j}(\beta)-\mathcal{F}_{\theta}\left(\beta^{+}|\mathcal{D}_{r},\mathcal{D}_{b},w\right)-\epsilon}{\sigma_{j}(\beta)}.$
@@
where $\mu_{j}(\beta)$ and $\sigma_{j}(\beta)$ represent the mean and standard deviation of the surrogate model at step $j$, $\beta^{+}$ denotes the current optimal observed point, $\epsilon$ is a tunable hyper-parameter. $\Phi(Z)$ and $\phi(Z)$ are the probability density function and cumulative density function of the Gaussian Process, respectively.
@@
Iterative Optimization Process. We start with an initial set of observations $\beta_{0},...,\beta_{h}$ by evaluating the objective function $\left\{\mathcal{F}\left(\beta_{i}|\mathcal{D}_{r},\mathcal{D}_{b},w,\theta\right)\right\}_{i=0}^{h}$ at a few selected points. Second, we use the observed data $\left(\beta_{j},\mathcal{F}\left(\beta_{j}|\mathcal{D}_{r},\mathcal{D}_{b},w,\theta\right)\right)$ to update the Gaussian Process model, refining the estimates of $\mu_{j}(\beta)$ and $\sigma_{j}(\beta)$. Third, we employ the Expected Improvement criterion to determine the next point $\beta_{j+1}$ and obtain $\left\{\mathcal{F}\left(\beta_{j+1}|\mathcal{D}_{r},\mathcal{D}_{b},w,\theta\right)\right\}_{i=0}^{h}$. We continue the process of updating the model and selecting new points until a stopping criterion is met. Finally, we select the point $\beta^{+}$ with the best observed value of the objective function $\mathcal{F}_{\theta}\left(\beta^{+}|\mathcal{D}_{r},\mathcal{D}_{b},w\right)$ as the optimal solution.
A.4 Experimental Setting
@@
Datasets We conduct experiments on various classification and generation tasks and examples of each dataset are shown in Tables 9 and 10. We collect all datasets from Huggingface. The train sets of these datasets are regarded as example datasets and the test datasets are used to evaluate the performance of ICL. We randomly subsample examples from the test dataset to verify the performance of the imbalanced annotated dataset. We generate categories for Natural question (Kwiatkowski et al., 2019) using ChatGPT-3.5-Turbo. The categories include people, time, geography, culture, and specialized knowledge.
@@
    Random selects demonstrations randomly from an example set without repetition (Min et al., 2022).
@@
    TopK retrieves demonstrations that are semantically similar to a test query sample (Liu et al., 2022).
@@
    DPP uses the original BERT embeddings as mentioned above without fine-tuning, and adopts MAP inference for subset retrieval (Ye et al., 2023a).
@@
    VoteK proposes an unsupervised and graph-based selective annotation method to select diverse and representative demonstrations (Su et al., 2023).
@@
    ConE searches for demonstrations by minimizing the difference in cross-entropy between the test input and the demonstrations (Peng et al., 2024).
@@
    ByCS assumes that an accurate inverse likelihood probability will lead to an accurate posterior probability and selects demonstrations based on their inverse inference results (Wang et al., 2024).
Experiment details We run our experiments on NVIDIA GeForce RTX 4090 and NVIDIA L40 GPU, and implement all methods by PyTorch and transformers. Our code is inspired by OpenICL (Wu et al., 2023). We thank the authors for releasing their code.
A.5 Inference
Perplexity For classification tasks, we compute the sentence perplexity for each sequence formed by concatenating the input with each candidate answer (Brown et al., 2020; Wu et al., 2023). Specifically, for each input instance $\mathrm{\mathbf{x}}$ is paired potential label set $\mathcal{Y}$, where $\mathcal{Y}$ represents the set of possible classes (e.g., Sports, Business, etc.). Then, for each possible label $\mathrm{\mathbf{y}}\in\mathcal{Y}$, we concatenate each tokenized input-output pair $(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})$, and obtain the corresponding tokenized sequence $\mathrm{\mathbf{c}}=(z_{1},...,z_{|\mathrm{\mathbf{c}}|})=(x_{1},...,x_{|\mathrm{\mathbf{x}}|},y_{1},...,y_{|\mathrm{\mathbf{y}}|})$, where $|\mathrm{\mathbf{c}}|=|\mathrm{\mathbf{x}}|+|\mathrm{\mathbf{y}}|$. Now, the perplexity of $\mathrm{\mathbf{c}}$ is calculated as:
@@ A.5 Inference
Direct For text generation tasks, we represent candidate answers using tokens from the vocabulary and select the final prediction based on the one with the highest probability (Brown et al., 2020; Wu et al., 2023). To reduce computational cost, we set the maximum new tokens to 50 and limit unnecessary token generation.
A.6 More empirical results
A.6.1 The effect of imbalanced dataset
In this section, we show the effect of imbalanced datasets on the performance of in-context learning with different model architectures and various numbers of subsets. Figures LABEL:figure:a1 (b) and (c) report that imbalanced datasets significant degrade ICL’s performance of Llama3-8B (AI@Meta, 2024) and Mistral-7B (Jiang et al., 2023). This illustrates that the negative effect of imbalanced datasets can be commonly observed in various large language models. Additionally, Figure LABEL:figure:a1 (a) also reports a larger set of demonstrations does not enhance the performance of ICL. Specifically, imbalanced datasets also negatively impact the performance of in-context learning.
A.6.2 The failure of classical rebalance methods
Figures LABEL:figure:a3 (a) and (b) reports the Kullback-Leibler (KL) divergence between $P_{c}(X|Y)$ and $P_{t}(X|Y)$ for each class in the Yelp and Yahoo datasets, respectively. Notably, the KL divergence between $P_{c}(X|Y)$ and $P_{t}(X|Y)$ for each class is unequal to zero and the assumption that $\beta$ = 0 is difficult to be satisfied.
A.6.3 Main results for each dataset
@@
A.6.4 Is our method effective with different model architectures?
In this section, we provide the detailed results of Llama3-8B (AI@Meta, 2024) and Mistral-7B (Jiang et al., 2023). Figures LABEL:figure:a4 (a) and (b) report the average accuracy of four classification tasks across six selection methods. We can clearly observe that imbalanced annotated datasets significantly hurt the performance of in-context learning, which is consistent with our previous findings. Meanwhile, Figures LABEL:figure:a4 (a) and (b) also demonstrate that integrating our method can improve the performance of ICL against imbalanced annotated datasets.
A.6.5 Can our method improve real-world imbalanced dataset?
We further verify the effectiveness of our method on a real-world class-imbalanced dataset (Emotion (Saravia et al., 2018)). .We select demonstrations from the training set and construct a balanced test set from the original test set. Figure 18 summarizes the test accuracy across six existing selection methods using baselines, classical rebalance methods, and our method on the Emotion dataset. In particular, the proposed method is able to consistently improve the existing selection methods and outperforms existing rebalance methods in test accuracy.
@@ A.6.5 Can our method improve real-world imbalanced dataset?
Imbalance Ratio 0 10 50 100 0 10 50 100
Random 12.33$\pm$0.62 15.2$\pm$1.45 18.53$\pm$1.32 20.13$\pm$0.98 13.33$\pm$0.19 18.53$\pm$1.05 20.07$\pm$0.09 20.33$\pm$0.34
+Ours 12.53$\pm$0.47 19.87$\pm$0.96 20.93$\pm$2.38 20.47$\pm$1.33 13.47$\pm$0.25 16.13$\pm$2.54 20.20$\pm$1.18 21.07$\pm$0.62
TopK 28.73$\pm$0.09 28.47$\pm$0.74 24.40$\pm$1.80 23.40$\pm$0.86 30.60$\pm$0.00 28.13$\pm$1.15 24.20$\pm$0.65 23.07$\pm$0.66
+Ours 28.80$\pm$0.00 29.47$\pm$0.50 27.87$\pm$0.96 27.20$\pm$0.49 30.40$\pm$0.00 31.00$\pm$1.30 29.13$\pm$0.93 28.33$\pm$0.90
DPP 29.80$\pm$0.16 27.13$\pm$0.68 25.47$\pm$1.27 24.13$\pm$0.5 32.47$\pm$0.09 28.13$\pm$0.47 24.00$\pm$0.43 23.00$\pm$0.75
+Ours 29.60$\pm$0.00 28.67$\pm$1.79 28.67$\pm$0.81 27.60$\pm$0.86 31.87$\pm$0.75 30.87$\pm$1.39 28.93$\pm$1.00 27.93$\pm$1.06
VoteK 28.67$\pm$0.09 25.67$\pm$0.81 22.73$\pm$1.23 22.93$\pm$0.41 31.57$\pm$0.29 27.67$\pm$0.82 22.60$\pm$0.65 21.93$\pm$0.25
+Ours 28.13$\pm$0.09 27.53$\pm$1.72 27.00$\pm$0.75 27.47$\pm$1.32 31.80$\pm$0.85 30.07$\pm$0.34 28.40$\pm$1.40 27.20$\pm$1.02
ConE 26.73$\pm$0.47 25.40$\pm$1.34 22.87$\pm$0.82 21.27$\pm$1.16 30.20$\pm$0.00 27.73$\pm$0.47 22.87$\pm$0.94 22.27$\pm$0.38
+Ours 27.40$\pm$0.82 26.73$\pm$1.27 27.60$\pm$1.30 27.27$\pm$1.16 30.27$\pm$0.34 29.07$\pm$1.20 29.13$\pm$0.93 26.80$\pm$0.49
ByDC 29.47$\pm$0.52 27.13$\pm$0.34 25.20$\pm$1.02 23.53$\pm$1.57 32.13$\pm$0.38 27.93$\pm$0.66 24.13$\pm$0.50 22.27$\pm$0.25
+Ours 29.67$\pm$0.50 28.80$\pm$1.42 29.20$\pm$0.71 27.13$\pm$1.05 31.53$\pm$1.23 30.07$\pm$1.31 28.60$\pm$1.14 27.8$\pm$0.75
Baselines 25.96$\pm$0.33 24.83$\pm$0.89 23.20$\pm$1.24 22.57$\pm$0.91 28.38$\pm$0.16 26.35$\pm$0.77 22.98$\pm$0.54 22.15$\pm$0.44
+Ours 26.02$\pm$0.31 26.85$\pm$1.28 26.88$\pm$1.15 26.19$\pm$1.04 28.26$\pm$0.57 27.87$\pm$1.35 27.40$\pm$1.10 26.52$\pm$0.81

### v2

@@ Appendix C Estimating Conditional Bias by Bayesian optimization
$\mathcal{F}_{\theta}\left(\bm{\beta}\right)=\mathcal{N}\left(\mu(\bm{\beta}),\sigma^{2}(\bm{\beta})\right),$ (12)
@@
where $\mathcal{N}$ is the Gaussian distribution with a mean function $\mu(\bm{\beta})$ and a covariance function $\sigma^{2}(\bm{\beta})$. For a data point $\bm{\beta}_{j}$ sampled from the Gaussian process $\mathcal{N}$, we compute its corresponding function values $\mathcal{F}_{\theta}\left(\bm{\beta}_{j}\right)$.
@@
Acquisition Function. We use the Expected Improvement (EI) criterion to select the point $\bm{\beta}_{j+1}$ with the maximum expected improvement as the next query point:
@@
$\operatorname{EI}(\bm{\beta})$ =$=\left\{\begin{array}[]{ll}\left(\mu_{j}(\bm{\beta})-\mathcal{F}_{\theta}\left(\bm{\beta}^{+}\right)-\epsilon\right)\Phi(Z)+\sigma_{j}(\bm{\beta})\phi(Z)&\text{ if }\sigma_{j}(\bm{\beta})>0,\\ 0&\text{ if }\sigma_{j}(\bm{\beta})=0.\end{array}\right.$ (15)
@@
$Z$ $=\frac{\mu_{j}(\bm{\beta})-\mathcal{F}_{\theta}\left(\bm{\beta}^{+}\right)-\epsilon}{\sigma_{j}(\bm{\beta})}.$ (16)
@@
where $\mu_{j}(\bm{\beta})$ and $\sigma_{j}(\bm{\beta})$ represent the mean and standard deviation of the surrogate model at step $j$, $\bm{\beta}^{+}$ denotes the current optimal observed point, $\epsilon$ is a tunable hyper-parameter. $\Phi(Z)$ and $\phi(Z)$ are the probability density function and cumulative density function of the Gaussian Process, respectively.
@@
Iterative Optimization Process. We start with an initial set of observations $\bm{\beta}_{0},...,\bm{\beta}_{h}$ by evaluating the objective function $\left\{\mathcal{F}\left(\bm{\beta}_{i}\right)\right\}_{i=0}^{h}$ at a few selected points. Second, we use the observed data $\left(\bm{\beta}_{j},\mathcal{F}\left(\bm{\beta}_{j}\right)\right)$ to update the Gaussian Process model, refining the estimates of $\mu_{j}(\bm{\beta})$ and $\sigma_{j}(\bm{\beta})$. Third, we employ the Expected Improvement criterion to determine the next point $\bm{\beta}_{j+1}$ and obtain $\left\{\mathcal{F}\left(\bm{\beta}_{j+1}\right)\right\}_{i=0}^{h}$. We continue the process of updating the model and selecting new points until a stopping criterion is met. Finally, we select the point $\bm{\beta}^{+}$ with the best observed value of the objective function $\mathcal{F}_{\theta}\left(\bm{\beta}^{+}\right)$ as the optimal solution.
Appendix D Experimental Setting
D.1 Datasets
@@
Datasets We conduct experiments on various classification and generation tasks and examples of each dataset are shown in Tables 12 and 13. We collect all datasets from Huggingface. The train sets of these datasets are regarded as example datasets and the test datasets are used to evaluate the performance of ICL. We randomly subsample examples from the test dataset to verify the performance of the imbalanced annotated dataset. We generate categories for Natural question [13] using ChatGPT-3.5-Turbo. The categories include people, time, geography, culture, and specialized knowledge.
D.2 Baselines
@@
    Random selects demonstrations randomly from an example set without repetition [30].
@@
    TopK retrieves demonstrations that are semantically similar to a test query sample [4].
@@
    DPP uses the original BERT embeddings as mentioned above without fine-tuning, and adopts MAP inference for subset retrieval [17].
@@
    VoteK proposes an unsupervised and graph-based selective annotation method to select diverse and representative demonstrations [31].
@@
    ConE searches for demonstrations by minimizing the difference in cross-entropy between the test input and the demonstrations [15].
@@
    ByCS assumes that an accurate inverse likelihood probability will lead to an accurate posterior probability and selects demonstrations based on their inverse inference results [6].
D.3 Inference
Perplexity For classification tasks, we compute the sentence perplexity for each sequence formed by concatenating the input with each candidate answer [1, 71]. Specifically, for each input instance $\mathrm{\mathbf{x}}$ is paired potential label set $\mathcal{Y}$, where $\mathcal{Y}$ represents the set of possible classes (e.g., Sports, Business, etc.). Then, for each possible label $\mathrm{\mathbf{y}}\in\mathcal{Y}$, we concatenate each tokenized input-output pair $(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})$, and obtain the corresponding tokenized sequence $\mathrm{\mathbf{c}}=(z_{1},...,z_{|\mathrm{\mathbf{c}}|})=(x_{1},...,x_{|\mathrm{\mathbf{x}}|},y_{1},...,y_{|\mathrm{\mathbf{y}}|})$, where $|\mathrm{\mathbf{c}}|=|\mathrm{\mathbf{x}}|+|\mathrm{\mathbf{y}}|$. Now, the perplexity of $\mathrm{\mathbf{c}}$ is calculated as:
@@ D.3 Inference
Direct For text generation tasks, we represent candidate answers using tokens from the vocabulary and select the final prediction based on the one with the highest probability [1, 71]. To reduce computational cost, we set the maximum new tokens to 50 and limit unnecessary token generation.
D.4 Experiment details
Experiment details We run our experiments on NVIDIA GeForce RTX 4090 and NVIDIA L40 GPU, and implement all methods by PyTorch and transformers. Our code is inspired by OpenICL [71]. We thank the authors for releasing their code.
Appendix E More empirical results
E.1 Performance on different model architectures and sizes
We conduct experiments on various sized LLMs (including open-weight models and APIs). The results (vanilla/ours) in the table 2 below present the same phenomenon as the main experiments in the manuscript: the ICL performance of LLMs get worse at larger imbalanced ratios, and our method can significantly improve the performance, especially at large imbalanced ratios (e.g., 100).
Table 2: Average accuracy (%) of Vanilla / +Ours methods on the AgNews and Yahoo datasets with different model architectures and sizes. Bold numbers are superior results.
Imbalance Ratio 1 10 50 100
Vanilla/+Ours
OPT-6.7B 68.71/68.63 62.90/67.54 57.35/64.92 54.31/62.07
OPT-13B 72.22/72.84 68.04/70.12 62.28/68.16 58.12/64.33
OPT-30B 76.34/76.60 72.32/74.10 67.74/71.90 64.30/70.08
LLAMA-3-8B 78.90/79.10 75.20/77.16 69.32/75.82 65.72/73.20
LLAMA-3-70B 86.50/86.43 83.07/84.80 79.86/83.13 77.60/81.22
ChatGPT-3.5-Turbo 82.45/82.40 79.47/81.08 77.88/80.20 76.32/79.74
Gemini-2.0-Flash 81.91/82.09 79.95/81.86 77.19/80.31 75.01/79.82
E.2 Can our method improve real-world imbalanced dataset
We verify the effectiveness of our method on a real-world imbalanced dataset. The Emotion [12] dataset has a long-tailed distribution (refer to Figure 5). Table 3 shows that our method consistently improves existing selection methods and outperforms existing rebalancing methods, including over-sampling [25], under-sampling [26], stratified sampling [27], and re-weighting [9]. For example, with OPT-6.7B [19], using our method boosts the average Macro-F1 of six selection methods from 32.12 to 36.69, a direct improvement of 4.57. These results verify that our method is effective in improving ICL’s performance in real-world imbalanced scenarios.
Figure 5: An example of imbalanced dataset. In Emotion [12], a few sentiments make a large contribution and data tend to show a long-tailed distribution. For example, “joy“ and “sadness“ are head classes, while most other classes, such as “love“ and “surprise“, are tail classes.
Table 3: Average Macro-F1 metric on Emotion dataset. Bold numbers are superior results.
Methods Vanilla +Over-sampling +Under-sampling +Re-Weighting +Stratified-Sampling +Ours
Accuracy 32.22 33.74 34.22 30.27 31.60 36.69
E.3 Reweighting with conditional bias vs. output calibration?
The ICL is closely related to the concept of label bias [1], where language models are biased toward certain answers during few-shot learning. The subsequent works [41, 58, 59, 60] delved deeper into this challenge and proposed to address it by calibrating the model’s output probabilities to compensate for this bias. One may also ask: can a similar effect be achieved by calibrating the output of ICL with imbalanced demonstrations? In this ablation, we compare our method with some calibration methods, including CC [56], DC [41] and LOOC [60].
Table 4 presents that although both the demonstration selection and calibration methods improve the performance of ICL against imbalanced annotations, the calibration methods obtain inferior performance compared to our proposed method. This is because the calibration methods [56, 41, 60] only ensure that ICL with imbalanced demonstrations generates consistent output with that of balanced demonstrations. However, some effective selection methods, such as TopK [4] and DPP [23], are more likely to select examples belonging to the same class as the test sample for demonstrations and conduct imbalanced demonstrations. For these selection methods, the calibration methods inevitably degrade the performance of ICL, whereas our method can preserve their effectiveness.
@@ E.2 Can our method improve real-world imbalanced dataset
Imbalance Ratios 1 10 50 100
Vanilla 84.45 75.19 67.78 64.02
+CC 78.38 77.29 75.75 74.42
+DD 79.63 78.67 76.71 75.58
+LOOC 75.71 73.00 74.25 73.88
+Ours 84.26 83.35 80.28 75.28
E.4 Discussion of imbalanced test set
We also conducted experiments on an imbalanced test set. The table below presents the average Macro-F1 metric on four datasets: AgNews, Amazon, Yelp, and Yahoo. The results in Table 5 demonstrate that our method can improve ICL performance on an imbalanced test set. For example, on the AgNews and Yahoo datasets with 100 imbalance ratio, our approach improves the test Macro-F1 of vanilla ICL from 45.42 to 53.85 – a 8.43 of direct improvement.
Table 5: Average Macro-F1 metric of Vanilla / +Ours methods on the AgNews and Yahoo datasets with imbalanced test set. Bold numbers are superior results.
Imbalance Ratios 1 10 50 100
Vanilla 64.32 58.11 54.09 45.42
+Ours 64.52 61.67 57.89 53.85
E.5 Results of extreme imbalanced ratios
[…]
Table 6: Average accuracy (%) of Vanilla / +Ours methods on the AgNews and Yahoo datasets with extreme imbalance ratios. Bold numbers are superior results.
Imbalance Ratios 10 100 500 1000
Vanilla 62.90 54.31 46.80 44.52
+Ours 67.54 62.07 58.34 56.33
Improvement 4.64 7.76 11.54 11.81
E.6 The computational cost of our method
In this section, we illustrate the high efficiency of our method from the following aspects. Firstly, our method estimates the conditional bias by accessing low-cost model outputs and does not require extensive hyperparameter tuning. The table 7 reports the average computational cost (in minutes) between our method and parameter-efficient fine-tuning (PEFT) [72] on the AgNews dataset across various imbalanced ratios. For example, with an imbalanced ratio of 100, one iteration of Bayesian optimization can be completed in 2.51 minutes using OPT-6.7B on a single NVIDIA L40 GPU.
Besides, the bayesian optimization employed in our method is well-suited for efficiently optimizing non-differentiable and black-box functions with a few iterations [69, 73, 74]. In fact, our method can achieve significant improvement with minimal computational cost using 30 iterations and a balanced subset of size 100. Our method is acceptable in practice and only model weights $w+\beta$ one time for the whole test dataset.
The detailed settings of PEFT are as follows: Learning Rate (1e-5); LR Scheduler Type (cosine); Batch Size (8); Number of Training Epochs (5); Number of Warmup Steps (100); Weight Decay (0.01); Lora Alpha (32); Lora Dropout (0.05).
Table 7: Comparison of average computational cost (minute) per iteration using RCB and PEFT on AgNews. Bold numbers are superior results.
Imbalance Ratios 1 10 50 100
PEFT 72.35 35.25 27.52 24.38
RCB 3.04 2.58 2.53 2.51
E.7 Main results for each dataset
[…]
@@ E.7 Main results for each dataset
Imbalance Ratio 1 10 50 100 1 10 50 100
Random 42.07$\pm$0.69 36.97$\pm$0.64 35.93$\pm$0.71 35.47$\pm$0.38 43.67$\pm$0.44 38.70$\pm$0.60 36.27$\pm$0.91 35.57$\pm$0.09
+Ours 41.63$\pm$0.89 41.67$\pm$1.96 38.90$\pm$1.00 40.17$\pm$0.98 43.00$\pm$0.60 43.60$\pm$1.47 42.93$\pm$0.56 43.33$\pm$1.24
TopK 46.50$\pm$0.00 42.43$\pm$1.02 37.17$\pm$1.36 37.13$\pm$1.31 47.60$\pm$0.00 43.20$\pm$0.27 38.07$\pm$1.29 37.17$\pm$1.36
+Ours 46.13$\pm$0.49 45.27$\pm$0.38 42.47$\pm$1.29 41.23$\pm$0.36 47.30$\pm$0.40 46.57$\pm$0.58 43.60$\pm$0.27 40.87$\pm$1.22
DPP 46.70$\pm$0.00 41.53$\pm$0.36 38.30$\pm$0.20 37.73$\pm$0.64 48.10$\pm$0.00 43.37$\pm$0.18 39.27$\pm$0.56 37.77$\pm$0.58
+Ours 46.30$\pm$0.27 43.63$\pm$1.11 41.23$\pm$1.96 39.77$\pm$0.56 47.97$\pm$0.09 45.47$\pm$0.44 42.13$\pm$0.96 40.23$\pm$0.44
VoteK 46.87$\pm$0.05 40.90$\pm$0.64 37.00$\pm$1.39 37.03$\pm$1.56 48.03$\pm$0.04 43.50$\pm$0.53 37.57$\pm$0.62 36.73$\pm$0.98
+Ours 46.67$\pm$0.16 43.23$\pm$1.49 42.50$\pm$1.00 40.57$\pm$0.64 47.93$\pm$0.16 44.83$\pm$0.58 43.17$\pm$0.69 40.73$\pm$0.64
ConE 46.13$\pm$0.38 40.27$\pm$0.66 36.60$\pm$0.86 35.67$\pm$1.67 48.10$\pm$0.07 42.77$\pm$0.16 37.87$\pm$0.76 36.33$\pm$0.91
+Ours 46.33$\pm$0.44 41.60$\pm$1.20 40.80$\pm$0.93 39.23$\pm$0.91 48.00$\pm$0.13 44.57$\pm$0.84 42.93$\pm$0.62 39.87$\pm$0.42
ByDC 46.73$\pm$0.09 41.60$\pm$1.40 37.47$\pm$0.93 36.57$\pm$1.10 48.23$\pm$0.09 43.43$\pm$0.11 38.60$\pm$1.00 37.43$\pm$0.71
+Ours 47.07$\pm$0.49 44.27$\pm$0.58 41.37$\pm$1.58 40.03$\pm$1.09 48.03$\pm$0.24 45.47$\pm$0.96 42.67$\pm$0.24 40.57$\pm$0.51
Vanilla 45.83$\pm$0.20 40.62$\pm$0.79 37.08$\pm$0.91 36.60$\pm$1.11 47.29$\pm$0.11 42.49$\pm$0.31 37.94$\pm$0.86 36.83$\pm$0.77
+Ours 45.69$\pm$0.46 43.28$\pm$1.12 41.21$\pm$1.29 40.17$\pm$0.76 47.04$\pm$0.27 45.08$\pm$0.81 42.91$\pm$0.56 40.93$\pm$0.75